OpenStudy (anonymous):

What are the solutions of the quadratic equation? x^2-9x+20=0

OpenStudy (wolf1728):

Use the quadratic formula x = [--9 +-sqrt(81 -4*1*20)] / 2 x= 9 + 1 / 2 x = 5 The other solution for x = x = [--9 -sqrt(81 -4*1*20)] / 2 x = (9-1)/2 x = 4